angela drove 20 miles in 25 minutes. how long do it take her to drive 1 mile ?

Answer by richwmiller(17219)

r*t=d
r=d/t
20/(25/60)
4/(5/12)=48 mph
48*t=1
t=1/48 hour
t=1/48*60=1.25 minutes